### v3

# The Addition Theorem for the Algebraic Entropy of Torsion Nilpotent Groups

Abstract.
The Addition Theorem for the algebraic entropy of group endomorphisms of torsion abelian groups was proved by Dikranjan, Goldsmith, Salce and Zanardo [4]. It was later extended by Shlossberg [10] to torsion nilpotent groups of class $2$. As our main result, we prove the Addition Theorem for endomorphisms of torsion nilpotent groups of arbitrary nilpotency class. As an application, we show that if $G$ is a torsion nilpotent group, then for every $\phi\in\operatorname{End}(G)$ either the entropy is infinite or $h(\phi)=\log(\alpha)$ for a positive integer $\alpha$. We further obtain, for automorphisms of locally finite groups, the Addition Theorem with respect to every term of the upper central series; in particular, it holds for automorphisms of $\omega$-hypercentral groups. Finally, we establish a reduction principle: if $\mathfrak{X}$ is a class of locally finite groups closed under taking subgroups and quotients, then the Addition Theorem for endomorphisms holds in $\mathfrak{X}$ if and only if it holds for locally finite groups generated by bounded sets.
Key words and phrases:
Addition Theorem, algebraic entropy, nilpotent groups, solvable groups, locally finite groups
2020 Mathematics Subject Classification:
28D20, 20F16, 20F18, 20F50
1. Introduction
Algebraic entropy is a numerical invariant measuring the asymptotic growth of trajectories of finite subsets under iterates of group endomorphisms. Since its introduction by Adler, Konheim and McAndrew [1], algebraic entropy has been widely studied, especially in the abelian setting, where a rich and well-understood theory has been developed; see, for example, [9, 11, 4, 3]. In particular, Weiss [11] and Peters [9] established Bridge Theorems connecting algebraic entropy with topological entropy.
A fundamental structural property of algebraic entropy is the Addition Theorem (AT), which describes how entropy behaves with respect to short exact sequences of groups that are invariant under a given endomorphism. Roughly speaking, the Addition Theorem expresses the entropy of an endomorphism as the sum of the entropies induced on a normal subgroup and on the corresponding quotient.
In the abelian setting, the Addition Theorem was first proved for torsion abelian groups by Dikranjan, Goldsmith, Salce and Zanardo [4], and later extended to all abelian groups by Dikranjan and Giordano Bruno [3] (see also [1, 9, 11] for classical origins and early developments). Nevertheless, Giordano Bruno and Spiga showed in [8] that the Addition Theorem fails in general for automorphisms of metabelian groups. Beyond the abelian case, the Addition Theorem has been established for certain classes of locally finite non-abelian groups. In particular, the case of torsion nilpotent groups of class $2$ was settled by Shlossberg [10]; see also [4, 7, 6, 12] for related results in the locally finite setting.
The purpose of this paper is to substantially extend the scope of the Addition Theorem in the non-abelian torsion setting. Our main result shows that the Addition Theorem holds for endomorphisms of torsion nilpotent groups of arbitrary nilpotency class. This provides a broad new class of non-abelian groups for which the behavior of algebraic entropy with respect to extensions is completely understood.
As a first application, we obtain a dichotomy for the algebraic entropy of endomorphisms of torsion nilpotent groups: if $G$ is torsion nilpotent and $\phi\in\operatorname{End}(G)$ has finite entropy, then $h(\phi)=\log(\alpha)$ for some $\alpha\in{\mathbb{N}_{+}}$.
We then turn to general locally finite groups. Using the nilpotent case as a key input, we show that for every locally finite group $G$ and every automorphism $\varphi\in\operatorname{Aut}(G)$, the Addition Theorem holds with respect to each term $Z_{n}(G)$ of the upper central series. As a consequence, in Section 5 we show that the Addition Theorem holds for automorphisms of $\omega$-hypercentral groups.
Finally, we establish in Section 5 a reduction principle for classes of locally finite groups closed under taking subgroups and quotients, showing that the Addition Theorem for endomorphisms holds in such a class if and only if it holds for locally finite groups generated by bounded sets.
2. Preliminaries
We now restrict our attention to locally finite groups. This allows us to work solely with finite subgroups instead of finite subsets. We can also replace the use of $\limsup$ by an actual limit, relying on the finite combinatorial structure of trajectories in the locally finite setting (cf. [2]).
Remark 2.1.
We will repeatedly use the following well-known fact: if $(a_{n})_{n\in{\mathbb{N}_{+}}}$ is a subadditive sequence of real numbers, that is,
[…]
then the limit
$\lim_{n\to\infty}\frac{a_{n}}{n}$
exists and equals $\inf_{n\in{\mathbb{N}_{+}}}\frac{a_{n}}{n}$. This is a classical result due to Fekete.
Definition 2.2.
Let $G$ be a locally finite group and let $\phi\colon G\to G$ be an endomorphism. For a finite subgroup $F\leq G$ and $n\in{\mathbb{N}_{+}}$, set
$T_{n}(\phi,F):=F\,\phi(F)\cdots\phi^{n-1}(F).$
Since $F$ is finite and $G$ is locally finite, each $T_{n}(\phi,F)$ is finite. The algebraic entropy of $\phi$ with respect to $F$ is then defined by the limit
$H(\phi,F):=\lim_{n\to\infty}\frac{1}{n}\log\bigl|T_{n}(\phi,F)\bigr|,$
which exists by subadditivity (see [2, Proposition 5.1.10]). The algebraic entropy of $\phi$ on $G$ is
$h(\phi):=\sup\{\,H(\phi,F):F\leq G\text{ finite subgroup}\,\}.$
2.1. Notation and terminology
The sets of non-negative reals, non-negative integers and positive natural numbers are denoted by ${\mathbb{R}}_{\geq 0}$, $\mathbb{N}$} and $\mathbb{N}_{+}$}, respectively. Throughout the paper, the notation $H\unlhd G$ indicates that $H$ is a normal subgroup of $G$. An element $x$ of a group $G$ is torsion if the subgroup $\langle x\rangle$ generated by $x$ is finite. Recall that a group $G$ is locally finite if every finitely generated subgroup of $G$ is finite. The derived series of a group $G$ is defined by
$G^{(0)}=G,\qquad G^{(i+1)}=[G^{(i)},G^{(i)}]\quad\text{for }i\geq 0.$
A group $G$ is called solvable if there exists $k\in{\mathbb{N}}$ such that $G^{(k)}=\{1\}$. In particular, $G$ is metabelian if $G^{\prime\prime}=G^{(2)}=\{1\}$. We also write $\gamma_{1}(G)=G$, $\gamma_{i+1}(G)=[\gamma_{i}(G),G]$ for the lower central series. In particular, $G^{\prime}=\gamma_{2}(G)$ and if $G$ is nilpotent of class at most $n$, then $\gamma_{n+1}(G)=\{1\}$ and $\gamma_{n}(G)\leq Z(G)$, where $Z(G)$ denotes the center of $G$.
A group $G$ is said to be locally solvable (respectively, locally nilpotent) if every finitely generated subgroup of $G$ is solvable (respectively, nilpotent). It is well-known that every torsion locally solvable group is locally finite (see [5, Proposition 1.1.5]).
We recall that the upper central series of a group $G$ is defined inductively by $Z_{0}(G)=\{1\}$ and
$Z_{n+1}(G)/Z_{n}(G)=Z\bigl(G/Z_{n}(G)\bigr)\qquad(n\in{\mathbb{N}}).$
Each $Z_{n}(G)$ is a characteristic subgroup of $G$. A group $G$ is said to be $\omega$-hypercentral if
$G=\bigcup_{n\in{\mathbb{N}}}Z_{n}(G).$
Recall that hypercentral groups (in the transfinite sense) are locally nilpotent by a classical result of Mal’cev (see, e.g., [5, p. 8]). In particular, every $\omega$-hypercentral group is locally nilpotent.
We denote by $\operatorname{End}(G)$ the set of all endomorphisms of $G$, and by $\operatorname{Aut}(G)$ its subset of automorphisms.
If $\phi\in\operatorname{End}(G)$, then a subgroup $H$ of $G$ is called $\phi$-invariant if $\phi(H)\subseteq H$, and $H$ is $\phi$-stable if $\phi\in\operatorname{Aut}(G)$ and $\phi(H)=H$. A subgroup $H$ of $G$ is characteristic if $H$ is $\phi$-invariant for every $\phi\in\operatorname{Aut}(G)$, and fully characteristic if the same holds for every $\phi\in\operatorname{End}(G)$. For every $n$ we let $G[n]$ be the fully characteristic subgroup of $G$ generated by the set $\{x\in G:x^{n}=e\}$. We say that a subset $F\subseteq G$ is bounded if the orders of the elements of $F$ are uniformly bounded.
3. Auxiliary results
We fix the following terminology for the Addition Theorem in a class of groups, which will be used throughout the paper
Definition 3.1.
Let $\mathfrak{X}$ be a class of groups closed under taking subgroups and quotients.
  1. (1)
    We say that $\operatorname{AT}(G,\phi,H)$ holds for a group $G\in\mathfrak{X}$, $\phi\in\operatorname{End}(G)$ and a $\phi$-invariant normal subgroup $H$ of $G$ if
    $h(\phi)=h(\phi\upharpoonright_{H})+h(\bar{\phi}),$
    where $\bar{\phi}=\bar{\phi}_{G/H}\in\operatorname{End}(G/H)$ is the map induced by $\phi.$
  2. (2)
    The Addition Theorem holds in $\mathfrak{X}$ for endomorphisms if $\operatorname{AT}(G,\phi,H)$ holds for every $G\in\mathfrak{X},\ \phi\in\operatorname{End}(G)$ and every $\phi$-invariant normal subgroup $H$ of $G.$
  3. (3)
    The Addition Theorem holds in $\mathfrak{X}$ for automorphisms if $\operatorname{AT}(G,\phi,H)$ holds for every $G\in\mathfrak{X},\ \phi\in\operatorname{Aut}(G)$ and every $\phi$-stable normal subgroup $H$ of $G.$
In the remainder of the paper, we will apply this terminology to the classes of torsion solvable groups, torsion nilpotent groups, and locally finite groups.
We begin with a basic invariance property of the algebraic entropy (see [2, Lemma 5.1.7]).
Lemma 3.2.
Let $G$ and $H$ be groups, $\phi\in\operatorname{End}(G)$ and $\psi\in\operatorname{End}(H)$. If there exists an isomorphism $\xi:G\to H$ such that $\psi=\xi\phi\xi^{-1}$, then
$h(\phi)=h(\psi).$
We shall also use the following reduction for solvable groups.
Lemma 3.3 (Glue lemma for solvable groups).
Let $G$ be a solvable group, $\phi\in\operatorname{End}(G)$, and let $H\unlhd G$ be a $\phi$-invariant normal subgroup. Assume that the following three Addition Theorem statements hold:
  1. (1)
    $\operatorname{AT}(G,\phi,G^{\prime})$,
  2. (2)
    $\operatorname{AT}(H,\phi\!\upharpoonright_{H},H^{\prime})$,
  3. (3)
    $\operatorname{AT}(G/H,\bar{\phi},(G/H)^{\prime})$,
where $\bar{\phi}$ is the endomorphism induced by $\phi$ on $G/H$. Then $\operatorname{AT}(G,\phi,H)$ holds.
Proof.
The proof follows the formal decomposition appearing in [12, Proposition 5.9], replacing applications of the Addition Theorem for abelian groups by the induction hypothesis on the solvability class.
Since $G$ is solvable, all groups appearing below (namely $G^{\prime}$, $H^{\prime}$, $(G/H)^{\prime}$, their intersections, and the corresponding quotients) have strictly smaller solvability class than $G$. Hence the induction hypothesis applies to each of them.
Using $\operatorname{AT}(G,\phi,G^{\prime})$, we decompose the entropy of $\phi$ along the extension $G^{\prime}\unlhd G$. Similarly, $\operatorname{AT}(H,\phi\!\upharpoonright_{H},H^{\prime})$ yields a decomposition of $h(\phi\!\upharpoonright_{H})$, and $\operatorname{AT}(G/H,\bar{\phi},(G/H)^{\prime})$ yields a decomposition of $h(\bar{\phi})$.
The natural identifications
$(G/H)^{\prime}=G^{\prime}H/H,\qquad(G/H)/(G^{\prime}H/H)\cong G/G^{\prime}H,$
together with invariance of algebraic entropy under conjugation, allow one to combine these three equalities exactly as in [12, Proposition 5.9]. This yields
$h(\phi)=h(\phi\!\upharpoonright_{H})+h(\bar{\phi}),$
that is, $\operatorname{AT}(G,\phi,H)$. ∎
Remark 3.4.
Lemma 3.3 generalizes [12, Proposition 5.9], where the group is assumed to be metabelian. In the present setting, abelian arguments are replaced by induction on the solvability class.
We next record a useful consequence of the Addition Theorem.
Proposition 3.5.
Let $\mathfrak{X}$ be a class of torsion solvable groups closed under taking subgroups and quotients. If the Addition Theorem holds in $\mathfrak{X}$ for endomorphisms, then for every $G\in\mathfrak{X}$ and every $\phi\in\operatorname{End}(G)$ either $h(\phi)=\infty$ or $h(\phi)=\log(\alpha)$ for some $\alpha\in{\mathbb{N}_{+}}.$
Proof.
In case $G$ is abelian (that is, solvable of class $1$) the assertion follows from [4, Proposition 1.3]. Now suppose that the assertion holds for groups from $\mathfrak{X}$ which are solvable of class less than $n$ and let $G\in\mathfrak{X}$ be of class $n$ and $\phi\in\operatorname{End}(G)$ with $h(\phi)<\infty$. By our assumption, the Addition Theorem holds in $\mathfrak{X}$, so
$h(\phi)=h(\phi\upharpoonright_{G^{\prime}})+h(\bar{\phi}),$
where $\bar{\phi}\in\operatorname{End}(G/G^{\prime})$ is the induced endomorphism. Since $G^{\prime}$ is abelian and $G/G^{\prime}$ is solvable of class less than $n$, the induction hypothesis implies that there exist $\alpha,\beta\in{\mathbb{N}_{+}}$ such that $h(\phi\upharpoonright_{G^{\prime}})=\log(\alpha)$ and $h(\bar{\phi})=\log(\beta)$. Therefore
$h(\phi)=\log(\alpha)+\log(\beta)=\log(\alpha\beta),$
and we are done. ∎
The next fact collects inequality from [10] that will be crucial in the nilpotent case.
Fact 3.6.
[10, Proposition 4.1] Let $G$ be a group and $\phi\in\operatorname{End}(G).$ If $N$ is a $\phi$-invariant central torsion subgroup of $G$, then
$h(\phi)\geq h(\phi\upharpoonright_{N})+h(\bar{\phi}_{G/N}).$
Finally, we recall the main result of [10], which we will use as a black box.
Fact 3.7.
[10, Theorem 4.6] If $G$ is a torsion nilpotent group of class $2$, $\phi\in\operatorname{End}(G)$ and $H$ is a $\phi$-invariant normal subgroup of $G$, then $\operatorname{AT}(G,\phi,H)$ holds.
4. Proving the main result
In this section we extend Fact 3.7 to all torsion nilpotent groups.
Lemma 4.1 (Subadditivity for central extensions).
Let $G$ be a locally finite group, let $N\leq Z(G)$ be a $\phi$-invariant subgroup, and let $\bar{\phi}$ be the endomorphism induced by $\phi$ on $G/N$. Then
$h(\phi)\leq h(\phi\!\upharpoonright_{N})+h(\bar{\phi}_{G/N}).$
Proof.
Let $F\leq G$ be a finite subgroup and let $\pi:G\to G/N$ be the quotient map. Set $Q=\pi(F)$ and note that $\pi(T_{n}(\phi,F))=T_{n}(\bar{\phi},Q)$ for every $n\geq 1$. For each $n\geq 1$ define the (finite) subgroup
$L_{n}=\langle\,Q,\bar{\phi}(Q),\ldots,\bar{\phi}^{n-1}(Q)\,\rangle\leq G/N.$
Since $G/N$ is locally finite, each $L_{n}$ is finite. Fix $n\geq 1$ and choose a set-theoretic section $s_{n}:L_{n}\to G$ with $\pi\circ s_{n}=\mathrm{id}_{L_{n}}$ and $s_{n}(1)=1$. Define the finite correction set
$C_{n}=\{\,s_{n}(x)s_{n}(y)(s_{n}(xy))^{-1}:x,y\in L_{n}\,\}\subseteq N.$
We next define a finite subgroup $K_{n}\leq N$ which will control all “error terms” arising in the decomposition of elements of $T_{n}(\phi,F)$. Since $F$ is finite and $G$ is locally finite, the set $T_{n}(\phi,F)$ is finite. For each $t\in T_{n}(\phi,F)$ fix a representation
$t=a_{0}a_{1}\cdots a_{n-1}\qquad\text{with }a_{i}\in\phi^{i}(F)\ (0\leq i<n).$
For each such $t$ and each $i$ put
$u_{i}(t):=a_{i}\,\bigl(s_{n}(\pi(a_{i}))\bigr)^{-1}\in N.$
Let
$U_{n}:=\{\,u_{i}(t):t\in T_{n}(\phi,F),\ 0\leq i<n\,\}\subseteq N.$
Then $U_{n}$ is finite, and we set
$K_{n}:=\langle\,F\cap N,\ C_{n},\ U_{n}\,\rangle\leq N.$
Since $G$ is locally finite and $K_{n}$ is finitely generated, $K_{n}$ is a finite subgroup of $N$.
We claim that
$|T_{n}(\phi,F)|\leq|T_{n}(\bar{\phi},Q)|\cdot|T_{n}(\phi\!\upharpoonright_{N},K_{n})|.$
Indeed, we have $\pi(T_{n}(\phi,F))=T_{n}(\bar{\phi},Q)$. For each $y\in T_{n}(\bar{\phi},Q)$ set
$A_{y}:=T_{n}(\phi,F)\cap\pi^{-1}(y).$
Then
$T_{n}(\phi,F)=\bigsqcup_{y\in T_{n}(\bar{\phi},Q)}A_{y},$
so
$|T_{n}(\phi,F)|\leq|T_{n}(\bar{\phi},Q)|\cdot\max_{y}|A_{y}|.$
Fix $y$ and choose $g_{y}\in T_{n}(\phi,F)$ with $\pi(g_{y})=y$. Then $A_{y}\subseteq T_{n}(\phi,F)\cap g_{y}N$, and left multiplication by $g_{y}^{-1}$ induces a bijection between $T_{n}(\phi,F)\cap g_{y}N$ and $g_{y}^{-1}T_{n}(\phi,F)\cap N$. Therefore
$|A_{y}|\leq|g_{y}^{-1}T_{n}(\phi,F)\cap N|.$
Let $t\in T_{n}(\phi,F)$, and consider the fixed representation $t=a_{0}\cdots a_{n-1}$ with $a_{i}\in\phi^{i}(F)$ chosen above. For each $i$ we have $\pi(a_{i})\in\bar{\phi}^{i}(Q)\subseteq L_{n}$, so
$a_{i}=u_{i}(t)\,s_{n}\bigl(\pi(a_{i})\bigr)\quad\text{with }u_{i}(t)\in U_{n}\subseteq K_{n}.$
Substituting these expressions into the product for $t$ and using that $N\leq Z(G)$, we may commute all elements $u_{i}(t)$ to the left and obtain
$t=\bigl(u_{0}(t)u_{1}(t)\cdots u_{n-1}(t)\bigr)\cdot\bigl(s_{n}(\pi(a_{0}))\,s_{n}(\pi(a_{1}))\cdots s_{n}(\pi(a_{n-1}))\bigr).$
Since $s_{n}$ is not a homomorphism, we repeatedly use the identity
$s_{n}(x)s_{n}(y)=c(x,y)\,s_{n}(xy),\qquad c(x,y)\in C_{n},$
to rewrite the product of lifts. Iterating this identity yields
$s_{n}(\pi(a_{0}))\cdots s_{n}(\pi(a_{n-1}))=c(t)\,s_{n}(\pi(t)),$
where $c(t)$ is a product of elements of $C_{n}$. In particular, $c(t)\in\langle C_{n}\rangle\subseteq K_{n}$. Combining the above expressions, we obtain
$t=\eta(t)\,s_{n}(\pi(t)),\qquad\eta(t):=\bigl(u_{0}(t)\cdots u_{n-1}(t)\bigr)\,c(t)\in K_{n}\subseteq N.$
Since $K_{n}\subseteq T_{n}(\phi\!\upharpoonright_{N},K_{n})$, we have $\eta(t)\in T_{n}(\phi\!\upharpoonright_{N},K_{n})$ for every $t\in T_{n}(\phi,F)$. Now fix $y\in T_{n}(\bar{\phi},Q)$ and let $g_{y}\in T_{n}(\phi,F)$ be as above with $\pi(g_{y})=y$. If $x\in A_{y}$, then $\pi(x)=\pi(g_{y})$, hence $\pi(g_{y}^{-1}x)=1$ and so $g_{y}^{-1}x\in N$. Using $(\ast)$ for both $x$ and $g_{y}$, we may write
$g_{y}=\eta(g_{y})\,s_{n}(\pi(g_{y}))\quad\text{and}\quad x=\eta(x)\,s_{n}(\pi(x)).$
Since $x\in A_{y}$, we have $\pi(x)=\pi(g_{y})=y$, and hence
$g_{y}=\eta(g_{y})\,s_{n}(y),\qquad x=\eta(x)\,s_{n}(y).$
Therefore
$g_{y}^{-1}x$ $=\bigl(\eta(g_{y})\,s_{n}(y)\bigr)^{-1}\,\bigl(\eta(x)\,s_{n}(y)\bigr)$
$=s_{n}(y)^{-1}\,\eta(g_{y})^{-1}\,\eta(x)\,s_{n}(y).$
Since $\eta(g_{y}),\eta(x)\in K_{n}\leq N$ and $N\leq Z(G)$, conjugation by $s_{n}(y)$ acts trivially on $N$, and thus
$g_{y}^{-1}x=\eta(g_{y})^{-1}\eta(x)\in K_{n}\subseteq T_{n}(\phi\!\upharpoonright_{N},K_{n}).$
We have shown that $g_{y}^{-1}A_{y}\subseteq T_{n}(\phi\!\upharpoonright_{N},K_{n})$, and hence
$|A_{y}|=|g_{y}^{-1}A_{y}|\leq|T_{n}(\phi\!\upharpoonright_{N},K_{n})|.$
Substituting this uniform bound into the fiber decomposition
$|T_{n}(\phi,F)|=\sum_{y\in T_{n}(\bar{\phi},Q)}|A_{y}|\leq|T_{n}(\bar{\phi},Q)|\cdot\max_{y}|A_{y}|$
yields $(\dagger)$.
Taking $\frac{1}{n}\log$ in $(\dagger)$ and letting $n\to\infty$, we obtain
$H(\phi,F)\leq H(\bar{\phi},Q)+\limsup_{n\to\infty}\frac{1}{n}\log\bigl|T_{n}(\phi\!\upharpoonright_{N},K_{n})\bigr|.$
Although the limit defining $H(\phi|_{N},K)$ exists for every fixed finite subgroup $K\leq N$, the finite subgroups $K_{n}$ depend on $n$. For this reason we estimate the sequence $\frac{1}{n}\log|T_{n}(\phi|_{N},K_{n})|$ using a $\limsup$.
Let
$F_{n}(\phi\!\upharpoonright_{N})=\sup\{\log|T_{n}(\phi|_{N},K)|:\ K\leq N\text{ finite}\}.$
Then $\{F_{n}(\phi\!\upharpoonright_{N})\}_{n\geq 1}$ is subadditive, hence $\lim_{n\to\infty}\frac{F_{n}(\phi|_{N})}{n}$ exists and equals $h(\phi|_{N})$. Since $K_{n}$ is finite, we have
$\frac{1}{n}\log|T_{n}(\phi|_{N},K_{n})|\leq\frac{F_{n}(\phi|_{N})}{n}$
for every $n$, and therefore
$\limsup_{n\to\infty}\frac{1}{n}\log|T_{n}(\phi|_{N},K_{n})|\leq h(\phi|_{N}).$
Taking the supremum over all finite $F\leq G$ gives the desired inequality. ∎
By Fact 3.6(1) and Lemma 4.1 we immediately have:
Corollary 4.2.
Let $G$ be a locally finite group and $\phi\in\operatorname{End}(G)$. If $N$ is a $\phi$-invariant central subgroup of $G,$ then
$h(\phi)=h(\phi\upharpoonright_{N})+h(\bar{\phi}_{G/N}).$
Proposition 4.3 (Addition Theorem for central extensions for automorphisms).
Let $G$ be a locally finite group and let $\varphi\in\operatorname{Aut}(G)$. Then
$h(\varphi)=h(\varphi\!\upharpoonright_{Z(G)})+h(\bar{\varphi}_{G/Z(G)}).$
In other words, $\operatorname{AT}(G,\varphi,Z(G))$ holds for every automorphism of every locally finite group.
Proof.
Since $\varphi$ is an automorphism, $\varphi(Z(G))=Z(G)$. The claim follows from Corollary 4.2. ∎
Let $G$ be a torsion nilpotent group of class $n$ and consider its fully characteristic central subgroup $\gamma_{n}(G).$ By Corollary 4.2 we obtain the following.
Corollary 4.4.
Let $G$ be a torsion nilpotent group of class $n$ and $\phi\in\operatorname{End}(G).$ Then
$h(\phi)=h(\phi\upharpoonright_{\gamma_{n}(G)})+h(\bar{\phi}_{G/\gamma_{n}(G)}).$
Now we are ready to prove our main result.
Theorem 4.5.
Let $G$ be a torsion nilpotent group, let $\phi\in\operatorname{End}(G)$ and let $H$ be a $\phi$-invariant normal subgroup of $G$. Then $\operatorname{AT}(G,\phi,H)$ holds.
Proof.
By Lemma 3.3, it suffices to prove that $\operatorname{AT}(G,\phi,G^{\prime})$ holds for every such $G$ and $\phi$. So from now on let $H=G^{\prime}$. We prove the assertion using induction on the nilpotency class of $G.$ If $G$ is torsion nilpotent of class $1$, then it is abelian and the Addition Theorem holds by [4]. Let $G$ be a torsion nilpotent group of class $n$. Applying Corollary 4.4 we obtain
(4.1) $h(\phi)=h(\phi\upharpoonright_{\gamma_{n}(G)})+h(\bar{\phi}_{G/\gamma_{n}(G)}).$
Claim 1.
$h(\phi\upharpoonright_{G^{\prime}})=h(\phi\upharpoonright_{\gamma_{n}(G)})+h(\bar{\phi}_{G^{\prime}/\gamma_{n}(G)}).$
Proof of Claim 1.
The subgroup $G^{\prime}$ is torsion nilpotent of class less than $n$ for which the Addition Theorem holds by the induction hypothesis. ∎
Claim 2.
$h(\bar{\phi}_{G/\gamma_{n}(G)})=h(\bar{\phi}_{G^{\prime}/\gamma_{n}(G)})+h(\bar{\phi}_{G/G^{\prime}}).$
Proof of Claim 2.
By our assumption, the quotient $G/\gamma_{n}(G)$ is torsion nilpotent of class n1n-1. The subgroup $G^{\prime}/\gamma_{n}(G)$ is a $\bar{\phi}$-invariant normal subgroup of $G/\gamma_{n}(G)$. By the induction hypothesis, the Addition Theorem holds for $\bar{\phi}_{G/\gamma_{n}(G)}$. In view of Lemma 3.2 we have
$h(\bar{\phi}_{G/\gamma_{n}(G)})=h(\bar{\phi}_{G^{\prime}/\gamma_{n}(G)})+h(\bar{\phi}_{G/G^{\prime}}),$
as required. ∎
Combining Claim 1 and Claim 2 with (4.1) we obtain
$h(\phi\upharpoonright_{G^{\prime}})+h(\bar{\phi}_{G/G^{\prime}})=h(\phi\upharpoonright_{\gamma_{n}(G)})+h(\bar{\phi}_{G^{\prime}/\gamma_{n}(G)})+h(\bar{\phi}_{G/G^{\prime}})=h(\phi\upharpoonright_{\gamma_{n}(G)})+h(\bar{\phi}_{G/\gamma_{n}(G)})=h(\phi).$
That is, $\operatorname{AT}(G,\phi,G^{\prime})$ holds and we completed the proof. ∎
It was proved by Giordano Bruno and Spiga [8] that inner automorphisms of arbitrary groups have the same algebraic entropy as the identity. In particular, inner automorphisms have zero algebraic entropy on locally finite groups, and hence on torsion nilpotent groups.
On the other hand, in the torsion nilpotent setting this fact can also be recovered directly from Theorem 4.5, by applying the Addition Theorem to the central extension $Z(G)\trianglelefteq G$ and arguing by induction on the nilpotency class.
Corollary 4.6.
Let $G$ be a torsion nilpotent group and let $g\in G$. Then the inner automorphism $\iota_{g}\colon G\to G$, $\iota_{g}(x)=g^{-1}xg$, has zero algebraic entropy.
Using Theorem 4.5 and Proposition 3.5 we deduce the following result.
Corollary 4.7.
If $G$ is a torsion nilpotent group, and $\phi\in\operatorname{End}(G)$ is of finite entropy, then $h(\phi)=\log(\alpha)$ for some $\alpha\in{\mathbb{N}_{+}}.$
As an application of the Addition Theorem for torsion nilpotent groups, we now extend Proposition 4.3 from the center to all terms of the upper central series.
Proposition 4.8 (Addition Theorem for the $n$-center of automorphisms).
Let $G$ be a locally finite group and let $\varphi\in\operatorname{Aut}(G)$. Then for every $n\in{\mathbb{N}_{+}}$,
$h(\varphi)=h(\varphi\!\upharpoonright_{Z_{n}(G)})+h(\bar{\varphi}_{G/Z_{n}(G)}).$
In other words, $\operatorname{AT}(G,\varphi,Z_{n}(G))$ holds for every $n\in{\mathbb{N}_{+}}$.
Proof.
We argue by induction on $n$. For $n=1$ the claim is exactly Proposition 4.3.
Assume that the claim holds for some $n\geq 1$, and set $\pi\colon G\to G/Z_{n}(G)$. Let $\bar{\varphi}$ be the automorphism induced by $\varphi$ on $G/Z_{n}(G)$. Since $G$ is locally finite, so is $G/Z_{n}(G)$; hence Proposition 4.3 applied to $(G/Z_{n}(G),\bar{\varphi})$ yields
$h(\bar{\varphi})=h\!\left(\bar{\varphi}\!\upharpoonright_{Z(G/Z_{n}(G))}\right)+h\!\left(\overline{\bar{\varphi}}_{(G/Z_{n}(G))/Z(G/Z_{n}(G))}\right).$
Using $Z(G/Z_{n}(G))=Z_{n+1}(G)/Z_{n}(G)$ and $(G/Z_{n}(G))/Z(G/Z_{n}(G))\cong G/Z_{n+1}(G)$, we obtain
$h(\bar{\varphi}_{G/Z_{n}(G)})=h(\widetilde{\varphi}_{Z_{n+1}(G)/Z_{n}(G)})+h(\bar{\varphi}_{G/Z_{n+1}(G)}),$
where $\widetilde{\varphi}$ denotes the automorphism induced by $\varphi$ on $Z_{n+1}(G)/Z_{n}(G)$.
On the other hand, $Z_{n+1}(G)$ is a torsion nilpotent group (of class at most $n+1$), and $Z_{n}(G)\unlhd Z_{n+1}(G)$ is $\varphi$-stable. Therefore, Theorem 4.5 applied to $\varphi\!\upharpoonright_{Z_{n+1}(G)}$ gives
$h(\varphi\!\upharpoonright_{Z_{n+1}(G)})=h(\varphi\!\upharpoonright_{Z_{n}(G)})+h(\widetilde{\varphi}_{Z_{n+1}(G)/Z_{n}(G)}).$
Finally, by the inductive hypothesis,
$h(\varphi)=h(\varphi\!\upharpoonright_{Z_{n}(G)})+h(\bar{\varphi}_{G/Z_{n}(G)}),$
and combining the previous identities yields
$h(\varphi)=h(\varphi\!\upharpoonright_{Z_{n+1}(G)})+h(\bar{\varphi}_{G/Z_{n+1}(G)}),$
which completes the induction. ∎
5. A reduction principle for locally finite groups
We shall need the following basic monotonicity result along directed unions (see [2, Proposition 5.1.10]).
Proposition 5.1.
Let $G$ be a locally finite group and let $\varphi\in\operatorname{End}(G)$. Let $\{G_{i}\}_{i\in I}$ be a directed family of $\varphi$-invariant subgroups of $G$ such that
$G=\bigcup_{i\in I}G_{i}.$
Then
$h(\varphi)=\sup_{i\in I}h\bigl(\varphi\!\upharpoonright_{G_{i}}\bigr).$
Proof.
This is a special case of [2, Proposition 5.1.10], applied to the directed system $\{G_{i}\}_{i\in I}$ (with bonding maps given by inclusions). ∎
Proposition 5.2.
Let $G$ be a locally finite group and let $\{G_{i}\}_{i\in I}$ be a chain of subgroups of $G$ such that $G=\bigcup_{i\in I}G_{i}$.
  1. (1)
    (Endomorphisms). Assume that each $G_{i}$ is fully characteristic in $G$ and that $\operatorname{AT}(G_{i})$ holds for endomorphisms for every $i\in I$. Then $\operatorname{AT}(G)$ holds for endomorphisms.
  2. (2)
    (Automorphisms). Assume that each $G_{i}$ is characteristic in $G$ and that $\operatorname{AT}(G_{i})$ holds for automorphisms for every $i\in I$. Then $\operatorname{AT}(G)$ holds for automorphisms.
Proof.
We prove 1. Let $\varphi\in\operatorname{End}(G)$ and let $H\unlhd G$ be a $\varphi$-invariant normal subgroup. Since each $G_{i}$ is fully characteristic, both $G_{i}$ and $H\cap G_{i}$ are $\varphi$-invariant and normal in $G_{i}$. As $\operatorname{AT}(G_{i})$ holds (for endomorphisms), we have for every $i\in I$:
$h(\varphi\!\upharpoonright_{G_{i}})=h(\varphi\!\upharpoonright_{H\cap G_{i}})+h(\overline{\varphi}\!\upharpoonright_{G_{i}/(H\cap G_{i})}).$
Taking suprema and using Proposition 5.1 for the unions
$H=\bigcup_{i\in I}(H\cap G_{i}),\qquad G/H=\bigcup_{i\in I}(G_{i}H)/H\cong\bigcup_{i\in I}G_{i}/(H\cap G_{i}),$
we obtain
$h(\varphi)=h(\varphi\!\upharpoonright_{H})+h(\overline{\varphi}),$
and hence $\operatorname{AT}(G)$ holds for endomorphisms.
The proof of 2 is analogous: one replaces $\varphi\in\operatorname{End}(G)$ by $\phi\in\operatorname{Aut}(G)$ and uses that characteristic subgroups are $\phi$-stable. ∎
Corollary 5.3.
Let $G$ be a locally finite group. If $\operatorname{AT}(G[n!])$ holds for endomorphisms for every $n\in{\mathbb{N}_{+}}$, then $\operatorname{AT}(G)$ holds for endomorphisms.
Proof.
For a locally finite group $G$, the family $\{G[n!]\}_{n\in{\mathbb{N}_{+}}}$ forms a chain of fully characteristic subgroups whose union is $G$, so the conclusion follows from Proposition 5.21. ∎
Proposition 5.4.
Let $G$ be a torsion $\omega$-hypercentral group and let $\phi\in\operatorname{Aut}(G)$. Then the Addition Theorem holds for $(G,\phi)$; that is, for every $\phi$-stable normal subgroup $H\trianglelefteq G$ one has
$h(\phi)=h(\phi\!\upharpoonright_{H})+h(\bar{\phi}_{G/H}).$
Moreover, either $h(\phi)=\infty$ or $h(\phi)=\log(\alpha)$ for some $\alpha\in{\mathbb{N}_{+}}.$
Proof.
Since $G$ is $\omega$-hypercentral, we have $G=\bigcup_{n\in{\mathbb{N}_{+}}}Z_{n}(G)$. Each $Z_{n}(G)$ is a characteristic (hence $\phi$-stable) torsion nilpotent subgroup of $G$. By Theorem 4.5, $\operatorname{AT}(Z_{n}(G))$ holds for endomorphisms, and hence also for automorphisms. Therefore Proposition 5.22 yields $\operatorname{AT}(G)$ for automorphisms.
For the entropy dichotomy, note that a torsion $\omega$-hypercentral group is locally nilpotent and hence locally finite. Applying Proposition 5.1 to the directed chain $\{Z_{n}(G)\}_{n\in{\mathbb{N}_{+}}}$ we obtain
$h(\phi)=\sup_{n\in{\mathbb{N}_{+}}}h\bigl(\phi\!\upharpoonright_{Z_{n}(G)}\bigr).$
By Corollary 4.7, for each $n$ either $h(\phi\!\upharpoonright_{Z_{n}(G)})=\infty$ or $h(\phi\!\upharpoonright_{Z_{n}(G)})=\log(\alpha_{n})$ for some $\alpha_{n}\in{\mathbb{N}_{+}}$. If $h(\phi)<\infty$, then all these entropies are finite, so
$h(\phi)=\sup_{n\in{\mathbb{N}_{+}}}\log(\alpha_{n})=\log(\alpha),$
where $\alpha=\sup_{n\in{\mathbb{N}_{+}}}\alpha_{n}\in{\mathbb{N}_{+}}$. ∎
We conclude with an example showing that the previous proposition applies to torsion $\omega$-hypercentral groups that are not nilpotent.
Example 5.5.
Let $p$ be a prime and let
$\mathrm{UT}_{\infty}(p)=\bigcup_{n\geq 1}\mathrm{UT}_{n}(p)$
be the group of finitary upper unitriangular matrices over $\mathbb{F}_{p}$. Then $\mathrm{UT}_{\infty}(p)$ is torsion and $\omega$-hypercentral, since every element lies in some nilpotent subgroup $\mathrm{UT}_{n}(p)$, but it is not nilpotent, as $\mathrm{UT}_{n}(p)$ has nilpotency class n1n-1. In particular, by the previous proposition, $\operatorname{AT}(\mathrm{UT}_{\infty}(p),\varphi,H)$ holds for every $\varphi\in\operatorname{Aut}(\mathrm{UT}_{\infty}(p))$ and every $\varphi$-stable normal subgroup $H$ of $\mathrm{UT}_{\infty}(p)$.